OpenStudy (anonymous):

How many terms of the arithmetic sequence 2,4,6,8, ... add up to 60,762?

Parth (parthkohli):

\( \color{Black}{\Rightarrow S = {n \over 2}(1st + last) }\)

OpenStudy (anonymous):

The sum of integers from 1 to n equals n(n+1)/2. The sequence given is 2 times the sum from 1 to n. So our equation is: 30,381 = n(n+1)/2

OpenStudy (anonymous):

not much new here, sorry

OpenStudy (anonymous):

How do I use that when I don't have n?

OpenStudy (lgbasallote):

first find a_n \[a_n = a_1 + (n+1)d\] \[a_n = 2 + (n-1)2\] \[a_n = 2 + 2n - 2\] \[a_n = 2n\] now use the formula for sum \[s_n = \frac{n(a_1 + a_n)}{2}\] \[60762 = \frac{n(2 + 2n)}{2}\] \[60762 = \frac{2n+2n^2}{2}\] \[60762 = n + n^2\] \[n^2 + n - 6072 = 0\] now use \[n = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (lgbasallote):

uhh that should be \[n^2 + n - 60762 = 0\] then use quadratic formula..you'll get your answer ;)

OpenStudy (lgbasallote):

do you get it @derptastic878 ?

OpenStudy (anonymous):

Yes, thanks. Just trying to work out a number so big, haha.

OpenStudy (lgbasallote):

hehe

OpenStudy (anonymous):

Also I put in the wrong number and didn't notice it was a perfect root. x_x